# EgoDTM: Towards 3D-Aware Egocentric Video-Language Pretraining

### v1

@@ Abstract
Egocentric video-language pretraining has significantly advanced video representation learning. Humans perceive and interact with a fully 3D world, developing spatial awareness that extends beyond text-based understanding. However, most previous works learn from 1D text or 2D visual cues, such as bounding boxes, which inherently lack 3D understanding. To bridge this gap, we introduce EgoDTM, an Egocentric Depth- and Text-aware Model, jointly trained through large-scale 3D-aware video pretraining and video-text contrastive learning. EgoDTM incorporates a lightweight 3D-aware decoder to efficiently learn 3D-awareness from pseudo depth maps generated by depth estimation models. To further facilitate 3D-aware video pretraining, we enrich the original brief captions with hand-object visual cues by organically combining several foundation models. Extensive experiments demonstrate EgoDTM’s superior performance across diverse downstream tasks, highlighting its superior 3D-aware visual understanding. Our code will be released at https://github.com/xuboshen/EgoDTM.
@@ 1 Introduction
Figure 1: Comparison of different egocentric pretraining paradigms. While previous paradigms focus on text-based [33, 47, 78] or 2D spatial region-aware learning [72], EgoDTM incorporates 3D spatial information to enhance video representations.
@@
The development of embodied AI capable of fulfilling diverse societal roles has long been a goal in artificial intelligence [23, 46]. A promising path to achieving this vision involves developing egocentric AI, which can comprehend human activities by analyzing large-scale egocentric videos captured from a first-person perspective using wearable devices. These videos provide rich insights into hand-object interactions (HOI) [18, 71, 10, 36], offering a window into how individuals interact with nearby objects through their hands. With the emergence of large-scale egocentric datasets [20], video-language pretraining [33, 78] has become a dominant paradigm for learning egocentric video representations, significantly improving performance on downstream tasks such as video-text retrieval [10, 60] and action recognition [59, 28].
@@
Humans possess an innate ability to perceive and reason about 3D spatial relationships, effortlessly perceiving relative distances and spatial arrangements from visual cues alone [61, 25, 58]. While there is no universal definition of a model’s 3D awareness, we define it as the ability to infer 3D information from 2D images. However, models pretrained solely on 2D frames and general-purpose text (e.g., CLIP [49]) often struggle to develop robust 3D awareness [17, 39, 52]. To equip egocentric models with such 3D awareness, we wonder: Can video-language models better understand egocentric contexts by incorporating 3D-aware perception?
@@
Achieving this goal presents two key challenges. First, 3D representations are diverse, and obtaining corresponding labels is often costly. Recent works [70, 76] explore using multi-view images for 3D reconstruction to distill 3D awareness into 2D models. However, in egocentric scenarios, depth maps are more practical and easier to obtain. Depth maps provide direct 3D distance information and help distinguish salient objects from the background, offering crucial cues for spatial understanding. Unfortunately, existing egocentric datasets with depth maps [36, 51, 26] remain limited in both scale and diversity for large-scale pretraining. Recent foundation models for depth estimation [68, 69, 5] demonstrate strong out-of-domain generalization, enabling us to construct a large-scale depth-augmented egocentric dataset at a low cost.
@@
Second, effective 3D-aware pretraining requires bridging the modality gap between depth and text. As illustrated in Figure 1, common video-language pretraining [33, 47, 78] primarily relies on textual supervision, thus avoiding this challenge. Recent works [72, 43] explore region-aware pretraining using non-pixel-level cues, such as text and sparse object bounding boxes, which face only minor modality gaps. However, unlike these non-pixel-level cues, regressing the pixel-level outputs (e.g., depth maps) requires fundamentally different capabilities compared to predicting texts or bounding boxes, as suggested by previous studies [27, 31]. Additionally, depth estimation typically requires high-resolution inputs and multi-scale features, whereas video-language models rely on large batch sizes for contrastive learning, making direct integration nontrivial. Thus, designing an effective learning approach and enriching textual supervision with spatial information are crucial for successful 3D-aware pretraining.
@@
To address these aforementioned challenges, we introduce EgoDTM, a novel 3D-aware egocentric video-language model that learns video representation from depth maps and spatially informed captions. In addition to dual transformer encoders for video-text alignment, EgoDTM adopts a 3D-aware module for video pretraining and a data construction pipeline to enrich captions with spatial information. To adapt the 3D-aware module for the video-language framework, we propose a unified visual representation with a lightweight depth decoder, supervised by depth predictions from foundation models [68]. Specifically, the lightweight depth decoder takes as input the visual representation from the last layer of the video encoder and predicts the composition of discrete adaptive bins [3, 4, 57, 32] to estimate low-resolution depth maps. Moreover, we adopt off-the-shelf visual foundation models [56, 54] to create high-quality HOI masks by first detecting HOIs, then selecting key frames, and finally tracking bidirectionally. To enrich the original captions with spatial information, we leverage a large language model (LLM) guided by the temporally consistent HOI masks. Through 3D-aware video-language pretraining, EgoDTM improves visual generalization on downstream tasks involving egocentric HOIs.
@@ 2 Related Works
Egocentric Video-Language Pretraining. Egocentric video understanding [20, 7, 45, 75, 40, 66, 10, 64, 67, 29, 65] typically involves human daily interaction between hands and objects from a first-person perspective. Inspired by visual-language pretraining paradigms in third-person datasets [2, 38, 49, 43, 72], EgoVLP [33] firstly proposes to conduct egocentric video-language pretraining on the large-scale egocentric dataset Ego4D [20], which aims to learn video representations from massive video-text data via contrastive learning. Since EgoVLP, similar paradigms [33, 47, 78, 1] have gained large success in EgoHOI understanding, For example, LaViLa [78] employs a video-conditioned GPT-2 [48] and a T5 [50] rephraser to generate text descriptions for egocentric videos to expand video-text data. Despite the progress, pretraining with text alone often lacks precision in target localization. In response, some studies have focused on developing region-aware representations for EgoVLMs. For example, HelpingHands [72] proposes learning from noisy hand-object detection results, while HENASY [43] ensembles an additional hierarchical encoder to learn HOI region-aware representation. However, these methods are limited to 2D reasoning and lack an understanding of real-world 3D context. In this work, we take a step towards 3D-aware egocentric video-language pretraining.
@@
Figure 2: EgoDTM learns 3D-aware representations from depth and text. Our dual encoders are constructed using only transformers [15, 63, 49] with flash attention [11]. During pretraining, we conduct (1) 3D-aware video pretraining: we design a lightweight 3D-aware decoder to predict depth using visual feature maps, supervised by a teacher foundation model [69]. The decoder contains a plain feature pyramid to get multi-scale features, a depth-aware transformer decoder to process depth queries with video features, and the heads to predict depth maps; (2) Spatial-aware textual enrichment: we enhance captions with spatial information by organically combining foundation models in the detect-track-generate pipeline. Different green markers denote inconsistency of HOI predictions; identical ones indicate consistency.
@@
3D-Awareness of 2D Vision Models. As suggested by research in developmental psychology and psychophysics, we humans are capable of understanding 3D information, like depth and orientations, from only 2D visual signals [61, 25, 58]. With the growing interest in embodied intelligence [23, 41] and personal AI assistants [46], the demand for 3D-aware 2D visual models has become increasingly important. Although there is no widely accepted consensus to define and build effective 3D-aware models, some studies have started to assess the 3D awareness of visual foundation models. For example, Probing3D [17] demonstrated that CLIP, which is pretrained exclusively with text, is significantly limited in predicting depth and 3D surface normals. EysWideShut[62] revealed that CLIP struggles with 3D-related tasks, such as recognizing object orientation and understanding compositional contexts. More recently, relevant works have made endeavors to enable visual foundation models (e.g., CLIP) with 3D awareness by fine-tuning on 3D-aware data. For instance, FiT3D [70] finetunes visual foundation models with 3D Gaussian features reconstructed from multi-view images. SpatialVLM [8] conducts visual instruction tuning on metric depth-aware QA pairs for multi-modal large language models. Different from these learning paradigms, we define 3D-awareness as the model’s latent ability to estimate depth and pretrain an egocentric video-language model with a lightweight 3D-aware decoder.
@@ 3.1 Video-Language Model Architecture
The video-language model follows a dual-encoder design [2], consisting of a video encoder and a text encoder.
Video and text encoders. Our video encoder employs a plain vision transformer [15] backbone to process video tokens. The video input $V\in\mathbb{R}^{F\times H\times W\times 3}$ is divided into $N=\frac{F}{F_{0}}\times\frac{H}{H_{0}}\times\frac{W}{W_{0}}$ non-overlapping cubes of dimension $F_{0}\times H_{0}\times W_{0}\times 3$, where $F$ is the number of frames, $H$ and $W$ denote height and width. These cubes are combined with positional embeddings and processed by the video encoder to produce a feature map $\mathbf{v}_{Z}\in\mathbb{R}^{N\times C}$. Note that a ${\rm[CLS]}$ token is added to represent the global video embedding $\mathbf{v}_{\rm cls}$. For the text encoder, captions $T$ are tokenized via a Byte Pair Encoding (BPE) tokenizer [55] and encoded with a transformer initialized from CLIP [49] to output the sentence embedding $\mathbf{t}_{\rm cls}$. Both encoders adopt the flash-attention [11] to mitigate the memory bottleneck of the attention mechanism.
Video-text alignment loss. We use the InfoNCE loss to align video and text embeddings. For simplicity, we define the video-to-text loss within a batch =$\mathcal{B}=\{(\mathbf{v}_{i},\mathbf{t}_{i}\}_{i=1}^{B}$ as follows:
@@
Plain feature pyramid. Depth estimation methods [44, 3, 53] often utilize multi-scale features from intermediate backbone layers. However, video-language pretraining typically uses only the final-layer representations. To reconcile these differences, we adopt a simplified FPN [30] that projects the video feature map $\mathbf{v}_{Z}$ into multi-scale outputs $\{\mathbf{v}_{Z_{1\times}},\mathbf{v}_{Z_{2\times}},\cdots,\mathbf{v}_{Z_{m\times}}\}$, where $\mathbf{v}_{Z_{m\times}}$ denotes the feature map of size $(N\times m^{2})\times C$, and the size of the predicted depth maps equals $N\times m^{2}$. Moreover, inspired by the adaptive bins [3, 4, 57, 32] in depth estimation, we segment the depth range into intervals and predict distributions at each pixel over the interval centers.
Depth-aware transformer decoder. This module processes multi-scale features to generate depth-aware representations capable of estimating distinct distance ranges. The transformer decoder follows standard architecture as [6, 9], transforming $L$ learnable depth queries $\mathbf{Q}\in\mathbb{R}^{L\times C}$ into depth-aware representations $\mathbf{Z}\in\mathbb{R}^{L\times C}$ using self-attention and cross-attention mechanisms. The cross attention processes $\mathbf{Q}$ as queries and concatenates $(\mathbf{v}_{Z_{1\times}},\dots,\mathbf{v}_{Z_{m-1\times}})$ as key-value pairs $(k,v)$. The learned depth queries estimate depths at different distances, as illustrated by varied gray colors in Figure 2.
Bin head. We partition the depth range of $[0,1]$ into $L$ intervals. The bin head $\mathcal{F}(\cdot):\mathbb{R}^{C}\rightarrow[0,1]$ processes depth-aware representations $\{\mathbf{z}_{i}\}_{i=1}^{L}$ to predict widths of depth interval bins $\{w_{i}\}_{i=0}^{L}$, satisfying $\sum_{i=0}^{L}w_{i}=1$ where we denote $w_{0}=0$. The $i$-th depth interval is defined by boundaries $[\sum_{i=1}^{i-1}w_{i},\sum_{i=1}^{i}w_{i}]$, with its center computed as:
@@
3D-aware pretraining loss. Given the lack of depth annotations in large-scale egocentric videos [20], we generate pseudo-depth labels using monocular depth estimation foundation models [53, 68, 69]. The pretraining loss is defined as:
@@
where $\tilde{D}$ and $D_{\rm gt}$ are the predicted and ground-truths inverse depth maps within the range $[0,1]$, respectively.

### v2

@@ Abstract
Egocentric video-language pretraining has significantly advanced video representation learning. Humans perceive and interact with a fully 3D world, developing spatial awareness that extends beyond text-based understanding. However, most previous works learn from 1D text or 2D visual cues, such as bounding boxes, which inherently lack 3D understanding. To bridge this gap, we introduce EgoDTM, an Egocentric Depth- and Text-aware Model, jointly trained through large-scale 3D-aware video pretraining and video-text contrastive learning. EgoDTM incorporates a lightweight 3D-aware decoder to efficiently learn 3D-awareness from pseudo depth maps generated by depth estimation models. To further facilitate 3D-aware video pretraining, we enrich the original brief captions with hand-object visual cues by organically combining several foundation models. Extensive experiments demonstrate EgoDTM’s superior performance across diverse downstream tasks, highlighting its superior 3D-aware visual understanding. Code: https://github.com/xuboshen/EgoDTM.
@@ 1 Introduction
Figure 1: Comparison of egocentric pretraining paradigms. While previous paradigms focus on text-based lin2022egovlp; pramanick2023egovlpv2; lavila or 2D spatial region-aware learning zhang2023helping, EgoDTM incorporates 3D spatial information to enhance video representations.
@@
The development of embodied AI capable of fulfilling diverse societal roles has long been a goal in artificial intelligence hu2023robotsurvey; plizzari2023outlook. A promising path to achieving this vision involves developing egocentric AI, which can comprehend human activities by analyzing large-scale egocentric videos captured from a first-person perspective using wearable devices. These videos provide rich insights into hand-object interactions (HOI) fan2023arctic; zhan2024oakink2; epic-kitchens-100; liu2022hoi4d, offering a window into how individuals interact with nearby objects through their hands. With the emergence of large-scale egocentric datasets grauman2022ego4d, video-language pretraining lin2022egovlp; lavila has become a dominant paradigm for learning egocentric video representations, significantly improving performance on downstream tasks such as video-text retrieval epic-kitchens-100; actor_observer and action recognition charades-ego; egtea.
@@
Humans possess an innate ability to perceive and reason about 3D spatial relationships, effortlessly perceiving relative distances and spatial arrangements from visual cues alone spelke2010beyond; koenderink1992surface; shepard1971mental. While there is no universal definition of a model’s 3D awareness, we define it as the ability to infer 3D information from 2D images. However, models pretrained solely on 2D frames and general-purpose text (e.g., CLIP clip) often struggle to develop robust 3D awareness cvpr-2024probing3D; nips-man2024lexicon3d; ramakrishnan2024doesspatialcognitionemerge. To equip egocentric models with such 3D awareness, we wonder: Can video-language models better understand egocentric contexts by incorporating 3D-aware perception?
@@
Achieving this goal presents two key challenges. First, 3D representations are diverse, and obtaining corresponding labels is often costly. Recent works yue2024fit3d; zhang2024condense explore using multi-view images for 3D reconstruction to distill 3D awareness into 2D models. However, in egocentric scenarios, depth maps are more practical and easier to obtain. Depth maps provide direct 3D distance information and help distinguish salient objects from the background, offering crucial cues for spatial understanding. Unfortunately, existing egocentric datasets with depth maps liu2022hoi4d; ragusa2021meccano; Kwon2021h2o remain limited in both scale and diversity for large-scale pretraining. Recent foundation models for depth estimation depthanything; depthanythingv2; iclr2025depthpro demonstrate strong out-of-domain generalization, enabling us to construct a large-scale depth-augmented egocentric dataset at a low cost.
@@
Second, effective 3D-aware pretraining requires bridging the modality gap between depth and text. As illustrated in Figure 1, common video-language pretraining lin2022egovlp; pramanick2023egovlpv2; lavila primarily relies on textual supervision, thus avoiding this challenge. Recent works zhang2023helping; phan2024henasy explore region-aware video-language pretraining using non-pixel-level cues, such as text and sparse object bounding boxes, which face only minor modality gaps. However, unlike these non-pixel-level cues, regressing the pixel-level outputs (e.g., depth maps) requires fundamentally different capabilities compared to predicting texts or bounding boxes, as suggested by previous studies li2023maskdino; li2023flip. Additionally, depth estimation typically requires high-resolution inputs and multi-scale features, whereas video-language models rely on large batch sizes for contrastive learning, making direct integration nontrivial. Thus, designing an effective learning approach and enriching textual supervision with spatial information are crucial for successful 3D-aware pretraining.
@@
To address these aforementioned challenges, we introduce EgoDTM, a novel 3D-aware egocentric video-language model that learns video representation from depth maps and spatially informed captions. In addition to dual transformer encoders for video-text alignment, EgoDTM adopts a 3D-aware module for video pretraining and a data construction pipeline to enrich captions with spatial information. To adapt the 3D-aware module for the video-language framework, we propose a unified visual representation with a lightweight depth decoder, supervised by depth predictions from foundation models depthanything. Specifically, the lightweight depth decoder takes as input the visual representation from the last layer of the video encoder and predicts the composition of discrete adaptive bins bhat2021adabins; bhat2022localbins; shao2023iebins; li2024binsformer to estimate low-resolution depth maps. Moreover, we adopt off-the-shelf visual foundation models hoidetector; sam2 to create high-quality HOI masks by first detecting HOIs, then selecting key frames, and finally tracking bidirectionally. To enrich the original captions with spatial information, we leverage a large language model (LLM) guided by the temporally consistent HOI masks. Through 3D-aware video-language pretraining, EgoDTM improves visual generalization on downstream tasks involving egocentric HOIs.
@@ 2 Related Works
Egocentric Video-Language Pretraining. Egocentric video understanding grauman2022ego4d; openset-nips2023; plizzari2023can; zhang2022fine; mangalam2024egoschema; xu2023egopca; epic-kitchens-100; xu2023pov; xue2023learning; li2021egoexo; xu2025do typically involves human daily interaction between hands and objects from a first-person perspective. Inspired by visual-language pretraining paradigms in third-person datasets bain2021frozen; luo2022clip4clip; clip; phan2024henasy; zhang2023helping, EgoVLP (lin2022egovlp) firstly proposes to conduct egocentric video-language pretraining on the large-scale egocentric dataset Ego4D grauman2022ego4d, which aims to learn video representations from massive video-text data via contrastive learning. Since EgoVLP, similar paradigms lin2022egovlp; pramanick2023egovlpv2; lavila; ashutosh2023hiervl have gained large success in EgoHOI understanding, For example, LaViLa lavila employs a video-conditioned GPT-2 (gpt2) and a T5 (t5) rephraser to generate text descriptions for egocentric videos to expand video-text data. Despite the progress, pretraining with text alone often lacks precision in target localization. In response, some studies have focused on developing region-aware representations for EgoVLMs. For example, HelpingHands zhang2023helping proposes learning from noisy hand-object detection results, while HENASY phan2024henasy ensembles an additional hierarchical encoder to learn HOI region-aware representation. However, these methods are limited to 2D reasoning and lack an understanding of real-world 3D context. In this work, we take a step towards 3D-aware egocentric video-language pretraining.
@@
Figure 2: EgoDTM learns 3D-aware representations from depth and text. Our dual encoders are constructed using only transformers dosovitskiy2020image; vaswani2017attention; clip with flash attention dao2022flashattention. During pretraining, we conduct (1) 3D-aware video pretraining: we design a lightweight 3D-aware decoder to predict depth using visual feature maps, supervised by a teacher foundation model depthanythingv2. The decoder contains a plain feature pyramid to get multi-scale features, a depth-aware transformer decoder to process depth queries with video features, and the heads to predict depth maps; (2) Spatial-aware textual enrichment: we enhance captions with spatial information by organically combining foundation models in the detect-track-generate pipeline. Different green markers denote inconsistency of HOI predictions; identical ones indicate consistency.
@@
3D-Awareness of 2D Vision Models. As suggested by research in developmental psychology and psychophysics, we humans are capable of understanding 3D information, like depth and orientations, from only 2D visual signals spelke2010beyond; koenderink1992surface; shepard1971mental. With the growing interest in embodied intelligence hu2023robotsurvey; mccarthy2024survey2 and personal AI assistants plizzari2023outlook, the demand for 3D-aware 2D visual models has become increasingly important. Although there is no widely accepted consensus to define and build effective 3D-aware models, some studies have started to assess the 3D awareness of visual foundation models. For example, Probing3D cvpr-2024probing3D demonstrated that CLIP, which is pretrained exclusively with text, is significantly limited in predicting depth and 3D surface normals. EysWideShuttong2024eyes revealed that CLIP struggles with 3D-related tasks, such as recognizing object orientation and understanding compositional contexts. More recently, relevant works have made endeavors to enable visual foundation models (e.g., CLIP) with 3D awareness by fine-tuning on 3D-aware data. For instance, FiT3D yue2024fit3d finetunes visual foundation models with 3D Gaussian features reconstructed from multi-view images. SpatialVLM chen2024spatialvlm conducts visual instruction tuning on metric depth-aware QA pairs for multi-modal large language models. Different from these learning paradigms, we define 3D-awareness as the model’s latent ability to estimate depth and pretrain an egocentric video-language model with a lightweight 3D-aware decoder.
@@ 3.1 Video-Language Model Architecture
The video-language model basically consists of a video encoder and a text encoder.
Video and text encoders. Our video encoder employs a plain vision transformer dosovitskiy2020image backbone to process video tokens. The video input $V\in\mathbb{R}^{F\times H\times W\times 3}$ is divided into $N=\frac{F}{F_{0}}\times\frac{H}{H_{0}}\times\frac{W}{W_{0}}$ non-overlapping cubes of dimension $F_{0}\times H_{0}\times W_{0}\times 3$, where $F$ is the number of frames, $H$ and $W$ denote height and width. These cubes are combined with positional embeddings and processed by the video encoder to produce a feature map $\mathbf{v}_{Z}\in\mathbb{R}^{N\times C}$. Note that a $\rm[CLS]$} token is added to represent the global video embedding $\mathbf{v}_{\rm cls}$. For the text encoder, captions $T$ are tokenized via a Byte Pair Encoding (BPE) tokenizer bpe and encoded with a transformer initialized from CLIP clip to output the sentence embedding $\mathbf{t}_{\rm cls}$. Both encoders adopt the flash-attention dao2022flashattention to mitigate the memory bottleneck of the attention mechanism.
Video-text alignment loss. We use the InfoNCE loss to align video and text embeddings. For simplicity, we define the video-to-text loss within a batch $\mathcal{B}=\{(\mathbf{v}_{i},\mathbf{t}_{i}\}_{i=1}^{B}$ as follows:
@@
Plain feature pyramid. Depth estimation methods piccinelli2024unidepth; bhat2021adabins; ranftl2020midas often utilize multi-scale features from intermediate backbone layers. However, video-language pretraining typically uses only the final-layer representations. To reconcile these differences, we adopt a simplified FPN li2022simplefpn that projects the video feature map $\mathbf{v}_{Z}$ into multi-scale outputs $\{\mathbf{v}_{Z_{1\times}},\mathbf{v}_{Z_{2\times}},\cdots,\mathbf{v}_{Z_{m\times}}\}$, where $\mathbf{v}_{Z_{m\times}}$ denotes the feature map of size $(N\times m^{2})\times C$, and the size of the predicted depth maps equals $N\times m^{2}$. Moreover, inspired by the adaptive bins bhat2021adabins; bhat2022localbins; shao2023iebins; li2024binsformer in depth estimation, we segment the depth range into intervals and predict distributions at each pixel over the interval centers.
Depth-aware transformer decoder. This module processes multi-scale features to generate depth-aware representations capable of estimating distinct distance ranges. The transformer decoder follows standard architecture as carion2020detr; cheng2022mask2former, transforming $L$ learnable depth queries $\mathbf{Q}\in\mathbb{R}^{L\times C}$ into depth-aware representations $\mathbf{Z}\in\mathbb{R}^{L\times C}$ using self-attention and cross-attention mechanisms. The cross attention processes $\mathbf{Q}$ as queries and concatenates $(\mathbf{v}_{Z_{1\times}},\mathbf{v}_{Z_{2\times}},\dots,\mathbf{v}_{Z_{m-1\times}})$ as key-value pairs $(k,v)$. The learned depth queries estimate depths at different distances, as illustrated by varied gray colors in Figure 2.
Bin head. We partition the depth range of $[0,1]$ into $L$ intervals. The bin head $\mathcal{F}(\cdot):\mathbb{R}^{C}\rightarrow[0,1]$ processes depth-aware representations $\{\mathbf{z}_{i}\}_{i=1}^{L}$ to predict widths of depth interval bins $\{w_{i}\}_{i=0}^{L}$, satisfying $\sum_{i=0}^{L}w_{i}=1$ where we denote $w_{0}=0$. The $i$-th depth interval is defined by boundaries $\sum_{i=1}^{i-1}w_{i},\sum_{i=1}^{i}w_{i}]$, with its center computed as:
@@
3D-aware pretraining loss. Given the lack of depth annotations in large-scale egocentric videos grauman2022ego4d, we generate pseudo-depth labels using monocular depth estimation foundation models ranftl2020midas; depthanything; depthanythingv2. The pretraining loss is defined as:
@@
where $\tilde{D}$ and $D_{\rm gt}$ are the predicted and ground-truths inverse depth maps within the range $[0,1]$.
Figure 3: Qualitative results of depth estimation from EgoDTM and DepthAnythingv2 depthanythingv2(DAv2) on datasets including in-domain but unseen Ego4D validation set grauman2022ego4d, out-of-domain and unseen data of EK100 epic-kitchens-100, MECCANO ragusa2021meccano, and H2O Kwon2021h2o. Note that DAv2 operates with a high resolution of 512p, while EgoDTM uses a lower resolution input of 224p and generates a depth map at a resolution of 56p. Despite the lower resolution input, EgoDTM demonstrates intuitive generalization across unseen egocentric datasets with diverse environments, illuminations, backgrounds, and varying HOI object sizes.